OpenStudy (anonymous):

derivatives I cant seem to find the f' of these P=.015v^3 v=13 mph answer is 7.6w/mph I tried .015(v+h)^3-.015(v)^3 but that doesnt seem to work, also d=.05v^2+v Find R.O.C. when v=49mph answer is 5.9 tried same process got 4.9 ?

OpenStudy (anonymous):

first one f' is 3v^2=3(13)^2=507*.015=7.605 is that right?

OpenStudy (anonymous):

0.045v^2 = dP/dv Hence substituting we get 7.605 ROC would be dd/dv which is0.1v+1, hence 5.9 I dont understand what the problem is really :/

OpenStudy (anonymous):

how did you get .1v+1

OpenStudy (anonymous):

or how did you get .045v^2 you see how I am trying to work it

OpenStudy (anonymous):

all i know if f(x+h)-f(x)/h

OpenStudy (anonymous):

alright. when you differentiate, this happens dy/dx (x^n) = nx^n-1 hence 3*0.015v^3-1 = 0.045v^2. Then substitute. Same goes for the other question. Also, dy/dx(x) = 1 since x^1, 1*x (1-1) = x^0 = 1

OpenStudy (anonymous):

so what ever th

OpenStudy (anonymous):

so whatever the power is just multiply it to the coefficient -1? not really following

OpenStudy (anonymous):

do you not use F(x+h)-f(x)/h ?

OpenStudy (anonymous):

Multiply the power with the variable, and subtract one from the power. If you don't know this then i advise redoing your calculus because this is the very first thing they teach in acalculus class. Sorry i dont know that formula i learnt calculus from a book and i did it by some formulas i saw

OpenStudy (anonymous):

that formula is the first thing they taught me Im in calc 1 right now

OpenStudy (anonymous):

thank you for the help

OpenStudy (anonymous):

i will work on it

OpenStudy (anonymous):

Ohh i assume you're at the very beginning. Apologies i don't know what the h stands for here. You'll be introduced to it real soon:)

OpenStudy (anonymous):

i believe i understand your method much easier thank you

OpenStudy (anonymous):

Haha you're welcome :) Yes its easier than dividing everything and splitting the function :)

OpenStudy (anonymous):

true

OpenStudy (anonymous):

quick q

OpenStudy (anonymous):

on the second problem since x^0 is 1 .05(49)+1=3.45 not 5.9

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

.1(49)+1=5.9

OpenStudy (anonymous):

Haha its alright :) All teh best :)

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

youre welcome :)